OpenStudy (anonymous):

factor out 2^k from (4^k -k-2)2^(k+1)

OpenStudy (anonymous):

\[2^{k+1}=2^k+2^1.\] since it is just saying factor out only 2^k you can just leave everything and times your other part of the equation by 2 \[2^k(2(4^k-k-2))\]

OpenStudy (anonymous):

you can multiply the 2 but be careful as you'd need parenthesis around 4^k

OpenStudy (anonymous):

I know you answered this correctly but you do have a pretty important typo in the first line, so to clarify, \[2^{k+1} = 2^k * 2^1\]